please help me solve

2b^ - 4ac = bc where a = -2, b = 4, c = 3

Answer by checkley71(8403)

2*4^(2)-4*-2*3=4*3
2*16+24=12
32+24=12
56=12 WHAT ARE YOU TRYING TO PROVE????